OpenStudy (theraggedydoctor):

Josh and Mike live 13 miles apart. Yesterday, Josh started to ride his bicycle toward Mike's house. A little later Mike started to ride his bicycle toward Josh's house. When they met, Josh had ridden for twice the length of time as Mike and at four-fifths of Mike's rate. How many miles had Mike ridden when they met?

OpenStudy (dan815):

twice the time, at 4/5th of mike's rate

OpenStudy (anonymous):

is it 5 miles?

OpenStudy (anonymous):

\[2*\frac{ 4 }{ 5 }d = \frac{ 8 }{ 5 }\] \[\frac{ 8 }{ 5 }d + d = \frac{ 13 }{ 5 }\] \[\frac{ 13 }{ 5 }d = 13\] d = 5

OpenStudy (dan815):

since he travelled twice the time at 4/5th of his rate you have this ratio of distance between them 8/5 : 1 8/5 d : d 8/5 d + d = 13 d=5